# LMU-ASC 21/25

### v1

@@ 5.4 Relation to the Deformation Problem
with further contributions of ghosts and auxiliary fields. Combining this with the quadratic contribution from (30) we have
@@ 7 Discussion
Here $ev$ is realized by the path integral, while $1_{\psi}$ and $1_{\bar{\psi}}$ give the projection on the Fock space. The deformation $Q(A)$ was previously motivated by demanding that nilpotency implies the non-linear Yang-Mills equation Dai:2008bh . Here, we derive it from the path integral as a result of picture changing. In the process, we also derive a non-linear space-time action directly from the world line theory, which was previously unavailable, primarily due to the absence of trace on the relevant operator algebra Grigoriev:2021bes . The path integral provides such a trace in a larger algebra, at the price of generating higher form fields. The latter feature is a world-line realization of the familiar fact in the world-sheet approach of string theory that the operator product expansion does not commute with level truncation.
This larger vertex algebra looks somewhat like a familiar form of the duality invariant Yang-Mills theory realized on the $\mathcal{N}=1$ world line Cremonini:2025eds . Indeed, one can embed the latter in $\mathcal{N}=2$ with a suitable projection in the ghost sector. Perhaps this can lead to some insight into what the non-projected $\mathcal{N}=2$ theory describes. Conversely, the presence of $\mathcal{N}=2$ world-line SUSY enabled us to project out higher form fields in the external state, which was not possible in $\mathcal{N}=1$ (akin to the inconsistency of level truncation in string theory). This was instrumental in order to recover the deformation problem described in section 2. A further non-trivial observation is that with our choice of $Y$, the middle vertex did not give rise to interactions with higher form fields either. This is crucial for the consistency of the deformation problem. One may note some similarity with the presence of enhanced global symmetries in certain string backgrounds (e.g. Sen:2015uoa ; Mattiello:2019gxc ; Maccaferri:2019ogq ) except that the decoupling holds for an arbitrary space-time profile. Perhaps this could help in relation to level truncation in string field theory.

### v2

@@ 5.4 Relation to the Deformation Problem
with further contributions of ghosts and auxiliary fields. For (5.4) to be well-defined, we need to allow for tensors with values in an associative algebra rather than a Lie algebra. For concreteness, we will assume all fields to take values in $\mathfrak{u}^{\mathbb{C}}(n)$ in what follows. The colour indices are thus implicit. Combining (5.4) with the quadratic contribution from (31) we have
@@ 7 Discussion
Here $ev$ is realized by the path integral, while $1_{\psi}$ and $1_{\bar{\psi}}$ give the projection on the Fock space. The deformation $Q(A)$ was previously motivated by demanding that nilpotency implies the non-linear Yang-Mills equation Dai:2008bh . Here, we derive it from the path integral as a result of picture changing. In the process, we also derive a non-linear space-time action directly from the world line theory, which was previously unavailable, primarily due to the absence of trace on the relevant operator algebra Grigoriev:2021bes . The path integral provides such a trace in a larger algebra, at the price of generating higher form fields. The latter feature is a world line realization of the familiar fact in the world-sheet approach of string theory that the operator product expansion does not commute with level truncation.
This larger vertex algebra looks somewhat like a familiar form of the duality invariant Yang-Mills theory realized on the $\mathcal{N}=1$ world line Cremonini:2025eds . Indeed, one can embed the latter in $\mathcal{N}=2$ with a suitable projection in the ghost sector. Perhaps this can lead to some insight into what the non-projected $\mathcal{N}=2$ theory describes. Conversely, the presence of $\mathcal{N}=2$ world line SUSY enabled us to project out higher form fields in the external state, which was not possible in $\mathcal{N}=1$ (akin to the inconsistency of level truncation in string theory). This was instrumental in order to recover the deformation problem described in section 2. A further non-trivial observation is that with our choice of $Y$, the middle vertex did not give rise to interactions with higher form fields either. This is crucial for the consistency of the deformation problem. One may note some similarity with the presence of enhanced global symmetries in certain string backgrounds (e.g. Sen:2015uoa ; Mattiello:2019gxc ; Maccaferri:2019ogq ) except that the decoupling holds for an arbitrary space-time profile. Perhaps this could help in relation to level truncation in string field theory.